A discrete random variable X has the following probability distribution. 2 -1 0.4 0 0.3 f(x) a Find P(X>0). Select one: 0.6 1please fast and clear
Answer #1

2 X -U flog) 0.4 0.3 posto a Ef(xi) =10 6.4+0.3ta= 1 a = 10.7 a=0.3 P(x20) = P(x=0) + P(X=2) = 0.3to.3 - 0.6 P(X>,0) = 0.6 i